# Skolem and Positivity Completeness of Ergodic Markov Chains

### 01 Dec 2023 v3

@@ 1 Introduction
Markov Chains, owing to their formulation as stochastic matrices, are syntactically a special class of Linear Dynamical Systems. However, it is their semantics as transformers of probability distributions that earns them widespread attention as a natural mathematical framework to describe probabilistic systems. Markov Decision Processes, for instance, extend Markov Chains and are a probabilistic decision-making model fundamental to verification and Reinforcement Learning.
Given their importance, there is naturally an extensive body of work on model checking Markov Chains: see Baier-Katoen for a comprehensive set of references. Most of the focus has been on the verification of linear- and branching-time properties of Markov Chains through solving systems of linear equations, or linear programs. An alternative approach Agrawal2015 ; logicprob ; KVAK10 ; KwonGul is to consider specifications on the state distribution at each time step, e.g., whether the probability of being in a given state at the $n^{th}$ step is at least $1/4$. Decidability in this setting is a lot more inaccessible: Agrawal2015 ; logicprob only present incomplete or approximate verification procedures, while KVAK10 ; KwonGul owe their model-checking procedures to additional mathematical assumptions. The following decision problems are representative of the fundamental difficulties this approach is fraught with [^1] For technical reasons discussed in §2.1 of the Preliminaries, deciding whether there exist infinitely many $n\in\mathbb{N}$ such that $m_{ij}^{(n)}=r$ is actually a tractable problem.:
@@
    There exist infinitely many $n\in\mathbb{N}$ such that $m_{ij}^{(n)}\sim r$.
@@
It is here that the syntactic nature of Markov Chains as Linear Dynamical Systems is brought to the fore. In order to address these problems, we turn our attention to the simplest primitive in reasoning about Linear Dynamical Systems: the Linear Recurrence Sequence (LRS). An LRS over $\mathbb{Q}$ of order $k$ is an infinite sequence $\langle u_{n}\rangle_{n=0}^{\infty}$ of rational numbers satisfying a recurrence relation
@@
for all $n\in\mathbb{N}$, where $a_{0},\dots,a_{k-1}\in\mathbb{Q}$. An LRS over $\mathbb{Q}$ of order $k$ is uniquely specified by $2k$ rational numbers: $a_{0},\dots,a_{k-1}$ for the recurrence relation, and $u_{0},\dots,u_{k-1}$ for the initial terms. The following are important open decision problems for LRS:
@@
It is not difficult to use the Cayley-Hamilton Theorem to show that the Markov Reachability problems reduce to the corresponding LRS Problems above. We need the converse reductions to establish the difficulty of the former.
Techniques to perform these reductions were known since the work of Turakainen turakainen , albeit in the context of probabilistic automata. In fact, as shown in (bell2022, , Lemma 3.1), it is possible to use Turakainen’s techniques to reduce the Positivity Problem for LRS of order $k$ to Markov Reachability Problem 2 for Markov Chains of order $\leq k+3$ by crucially assuming that the initial terms $u_{0},\dots,u_{k-1}$ are non-negative. This assumption is not valid for instances of Skolem and Ultimate Positivity: while reducing these problems, the construction suffers from a blow-up in the order of the resulting Markov Chain.
The folkloric results are sharpened in oldresult : it is shown that the LRS Problems for order $k$ LRS reduce to Markov Reachability Problems for reducible and periodic Markov Chains of order $54k+$ 5. Number-theoretic hardness is formally shown for the Positivity and Ultimate Positivity Problems of order $6$ LRS in joeljames3 : decision procedures would enable one to quantify how well the continued fraction expansion of a given irrational number converges, an endeavour inaccessible to the means of contemporary number theory.
@@
    The current reduction demonstrates the hardness of verifying Markov Chains that are reducible and periodic, rather than ergodic (irreducible and aperiodic) Markov Chains the community most often considers.
@@
    Breakthroughs in the verification of Linear Dynamical Systems often result from restrictions on the dimensions skolem ; skolemv ; joeljames3 ; ouaknine2014positivity or spectrum ouaknine2014ultimate ; ouaknine2014positivity ; prefixindependent ; powerpositivity of the underlying matrix. A strong reduction needs to show that hardness persists despite the additional spectral properties (see, e.g. Theorem 2) enjoyed by ergodic Markov Chains, and also be faithful with respect to order while mapping LRS to stochastic matrices.
Our main contribution comprehensively addresses the above.
@@ 2.1 Linear Recurrence Sequences
It is straightforward to check that an LRS, as defined in the Introduction, can be computed through a companion matrix $\mathbf{A}$:
The characteristic polynomial of an LRS is given by $X^{k}-a_{k-1}X^{k-1}-\dots-a_{0}$; its roots are called the characteristic roots. The LRS is said to be non-degenerate if the ratio of two distinct characteristic roots is never a root of unity, and simple if the characteristic polynomial has no repeated roots.
The celebrated Skolem-Mahler-Lech Theorem skolem-sml ; mahler-sml ; lech-sml states that the set $\{n:u_{n}=0\}$ is the union of a finite set $Z$ and finitely many arithmetic progressions $c_{1}+b_{1}\mathbb{N},\dots,c_{d}+b_{d}\mathbb{N}$, where $0\leq c_{i}<b_{i}$ for all $i$. The values of $d,c_{i},b_{i}$ can all be effectively computed, whereas determining whether $Z$ is nonempty is precisely equivalent to the Skolem Problem. Berstel and Mignotte berstelmignotte showed that for any LRS $\langle u_{n}\rangle_{n=0}^{\infty}$, one can compute $B\geq 1$ such that for all $0\leq c<B$, the subsequence $\langle u_{nB+c}\rangle_{n=0}^{\infty}$ is a non-degenerate LRS which, as a consequence of the Skolem-Mahler-Lech Theorem, is either identically zero or has finitely many zeroes. It thus suffices to study non-degenerate LRS to solve the Skolem, Positivity and Ultimate Positivity Problems.
The Skolem Problem is known to be decidable for LRS of order up to $4$, see skolem ; skolemv . Very recently, there have been conditional decidability results for LRS of order $5$ bilu2022skolem . The Positivity Problem and Ultimate Positivity Problems are decidable up to order $5$, decidability at order $6$ would entail significant number-theoretic breakthroughs joeljames3 as mentioned in the Introduction. If we restrict ourselves to the class of simple LRS, then Ultimate Positivity is decidable ouaknine2014ultimate , and Positivity is decidable up to order $9$ ouaknine2014positivity . Neither decidability nor hardness results are known for the Positivity Problem for simple LRS of order $10$ and above.
@@ 3 The Reduction
We then observe that for all $n\geq 1$, $\mathbf{S}^{n}=\mathbf{S}$ and $\mathbf{M}^{n}=(\mathbf{S}+\mathbf{D})^{n}=\mathbf{S}+\mathbf{D}^{n}$, since $\mathbf{DS}=\mathbf{SD}=\mathbf{O}$. We let $r=s_{ij}$, and complete the reduction as follows (note that we actually reduce the complements of the Positivity and Ultimate Positivity Problems):
@@
It only remains to construct $\mathbf{D}$ and choose indices $i,j$, and constants $\eta,\rho$. We do so in a top-down fashion.

### 02 Feb 2024 v4

@@ 1 Introduction
Markov Chains, owing to their formulation in terms of stochastic matrices, are syntactically a special class of Linear Dynamical Systems. However, it is their semantics as transformers of probability distributions that earns them widespread attention as a natural mathematical framework to describe probabilistic systems. Markov Decision Processes, for instance, extend Markov Chains and are a probabilistic decision-making model fundamental to verification (Baier-Katoen, , Chapter 10) and Reinforcement Learning (mdp-rl, , Section 3).
Given their importance, there is naturally an extensive body of work on model checking Markov Chains: see Baier-Katoen for a comprehensive set of references. Most of the focus has been on the verification of linear- and branching-time properties of Markov Chains through solving systems of linear equations, or linear programs. An alternative approach Agrawal2015 ; logicprob ; KVAK10 ; KwonGul is to consider specifications on the state distribution at each time step, e.g., whether the probability of being in a given state at the $n^{th}$ step is at least $1/4$. Decidability in this setting is a lot more inaccessible: Agrawal2015 ; logicprob only present incomplete or approximate verification procedures, while KVAK10 ; KwonGul owe their model-checking procedures to additional mathematical assumptions. The following decision problems are representative of the fundamental difficulties this approach is fraught with:
@@
    There exist infinitely many $n\in\mathbb{N}$ such that $m_{ij}^{(n)}\sim r$ [^2] For technical reasons discussed in §2.1 of the Preliminaries, deciding whether there exist infinitely many $n\in\mathbb{N}$ such that $m_{ij}^{(n)}=r$ is actually a tractable problem.
@@
It is here that the syntactic nature of Markov Chains as Linear Dynamical Systems is brought to the fore. The essential tool we use to argue the difficulty of these problems is the Linear Recurrence Sequence (LRS). An LRS over $\mathbb{Q}$ of order $k$ is an infinite sequence $\langle u_{n}\rangle_{n=0}^{\infty}$ of rational numbers satisfying a recurrence relation
@@
for all $n\in\mathbb{N}$, where $a_{0}\neq 0,a_{1},\dots,a_{k-1}\in\mathbb{Q}$. An LRS over $\mathbb{Q}$ of order $k$ is uniquely specified by $2k$ rational numbers: $a_{0},\dots,a_{k-1}$ for the recurrence relation, and $u_{0},\dots,u_{k-1}$ for the initial terms. An LRS can be computed through its companion matrix $\mathbf{A}$:
The behaviour of an LRS is governed by the eigenvalues of $\mathbf{A}$, or equivalently, the roots of its characteristic polynomial $X^{k}-a_{k-1}X^{k-1}-\dots-a_{0}$. The LRS is said to be non-degenerate if the ratio of two distinct characteristic roots is never a root of unity, and simple if the characteristic polynomial has no repeated roots.
The following are key number-theoretic open decision problems for LRS:
@@
Number-theoretic hardness is formally shown for the Positivity and Ultimate Positivity Problems of order $6$ LRS in joeljames3 : decision procedures would enable one to quantify how well the continued fraction expansion of a given irrational number converges, an endeavour inaccessible to the means of contemporary number theory.
On the other hand, decidability results have been obtained through suitable restrictions. The Skolem Problem is known to be decidable for LRS of order up to $4$, see skolem ; skolemv . Very recently, there have been conditional decidability results for LRS of order $5$, and simple LRS bilu2022skolem . The Positivity Problem and Ultimate Positivity Problems are decidable up to order $5$ joeljames3 . If we restrict ourselves to simple LRS, then Ultimate Positivity is decidable ouaknine2014ultimate , and Positivity is decidable up to order $9$ ouaknine2014positivity . Neither decidability nor hardness results are known for the Positivity Problem for simple LRS of order $10$ and above.
Returning to our stochastic setting, we note that techniques to reduce the above LRS problems to Markov Reachability Problems were known since the work of Turakainen turakainen , albeit in the context of probabilistic automata [^3] It is not difficult to use the Cayley-Hamilton Theorem to show the reverse reduction, i.e. that the Markov Reachability problems reduce to the corresponding LRS Problems.. The folkloric results are sharpened in oldresult : it is shown that the LRS Problems for order $k$ LRS reduce to Markov Reachability Problems for reducible and periodic Markov Chains of order $54k+$ 5.
@@
    The best known reduction demonstrates the hardness of verifying Markov Chains that are reducible and periodic, rather than ergodic (irreducible and aperiodic) Markov Chains that the community most often considers.
@@
    Breakthroughs in the verification of Linear Dynamical Systems often result from restrictions on the dimensions joeljames3 ; skolem ; skolemv ; ouaknine2014positivity or spectrum ouaknine2014ultimate ; ouaknine2014positivity ; prefixindependent ; powerpositivity of the underlying matrix. A strong reduction needs to show that hardness persists despite the additional spectral properties (see, e.g. Theorem 2) enjoyed by ergodic Markov Chains, and also be faithful with respect to order while mapping LRS to stochastic matrices.
Our main result comprehensively addresses this issue.
@@ 2.1 Linear Recurrence Sequences
The celebrated Skolem-Mahler-Lech Theorem skolem-sml ; mahler-sml ; lech-sml states that the set $\{n:u_{n}=0\}$ is the union of a finite set $Z$ and finitely many arithmetic progressions $c_{1}+b_{1}\mathbb{N},\dots,c_{d}+b_{d}\mathbb{N}$, where $0\leq c_{i}<b_{i}$ for all $i$. The values of $d,c_{i},b_{i}$ can all be effectively computed, whereas determining whether $Z$ is nonempty is precisely the difficulty of the Skolem Problem. It is well known berstelmignotte ; everest that for any LRS $\langle u_{n}\rangle_{n=0}^{\infty}$, one can compute $B\geq 1$ such that for all $0\leq c<B$, the subsequence $\langle u_{nB+c}\rangle_{n=0}^{\infty}$ is a non-degenerate LRS which, by an observation of Lech lech-sml , is either identically zero or has finitely many zeroes. It thus suffices to study non-degenerate LRS to solve the Skolem, Positivity and Ultimate Positivity Problems.
@@ 3 The Reduction
We then observe that for all $n\geq 1$, $\mathbf{S}^{n}=\mathbf{S}$ and $\mathbf{M}^{n}=(\mathbf{S}+\mathbf{D})^{n}=\mathbf{S}+\mathbf{D}^{n}$, since $\mathbf{DS}=\mathbf{SD}=\mathbf{O}$. The property that $\mathbf{DS}=\mathbf{O}$ also gives us that $\mathbf{s}$ is the stationary distribution of $\mathbf{M}$: indeed, $\mathbf{MS}=\mathbf{S}^{2}+\mathbf{DS}=\mathbf{S}$. We let $r=s_{ij}$, and complete the reduction as follows (note that we actually reduce the complements of the Positivity and Ultimate Positivity Problems):
@@
Since $m_{ij}^{(n)}=s_{ij}+d_{ij}^{(n)}=r+d_{ij}^{(n)}$, it only remains to construct $\mathbf{D}$ and choose indices $i,j$, and constants $\eta,\rho$. We do so in a top-down fashion.